OpenStudy (anonymous):

F+5=5+f A distributive B communnitive C associative D identity

OpenStudy (kewlgeek555):

@Emily_13 : \(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\) Since OpenStudy does not allow you to giv direct answers, let me give you a hint :) A. Distributive Property Example:  B. Communative Property Example: 5 + 2 = 2 + 5 C. Associative Property Example: (2 + 7) + 5 = 2 + ( 7 + 5) D. Identity Property Example: 3 + 0 = 3; 9 + 0 = 9

OpenStudy (kewlgeek555):

For the Distributive Property Example, in the drawing, those are arrows. cx I don't draw very nice on the computer. cx

OpenStudy (kewlgeek555):

So, based on the examples I gave you, what do you think the answer is?

OpenStudy (anonymous):

B

OpenStudy (kewlgeek555):

Correct! f + 5 = 5 + f is an example of the Commutative Property! This shows that in an addition or multiplication equation, no matter what order you put it, it will be the same answer... 5 + 6 = 11 6 + 5 = 11 \[9 \times 8 =72\]\[8 \times 9 = 72\]

OpenStudy (anonymous):

Wow i thaught it was wrong

OpenStudy (kewlgeek555):

Well you were right. ;)